# Expected Shortfall LASSO

### v1

@@ 1 Introduction
We use a penalty function that accounts for differences in the variance of the regressors which has not been studied in the literature before for least squares estimators, to our knowledge, and which is borrowed from the penalized quantile regression literature, see Belloni and Chernozhukov, (2011). Although accounting for the variance is an appealing property on its own, we also prefer this penalty function as it means the penalization function of the coefficients in the VaR and ES estimation steps is equivalent. Indeed, we rely on the quantile regression estimator of Belloni et al., (2023) (from hereon BCMPW) to obtain the VaR predictions in the first step, although our results apply to any quantile estimator that satisfies the conditions.
As financial data is often dependent over time and heavy-tailed, we derive our results under conditions that allow for these properties. Specifically, we allow the data to be $\beta$-mixing sequences with finite moments of a certain order. To obtain our results, we develop a Fuk-Nagaev inequality for $\beta$-mixing sequences, which may be of independent interest. Using the Fuk-Nagaev inequality we obtain rates on the parameters and penalty parameter that are directly comparable to those for the penalized quantile regression estimator in BCMPW.
@@ 2.1 The model
where $\{U_{t}\}$ is a sequence of random errors satisfying $U_{t}|X_{t}\sim\text{Unif}(0,1)$, $\{X_{t}\}$ is a sequence of random vectors satisfying $X_{t}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, and $\alpha^{0}(\cdot)$ is some $p$-dimensional vector functional on $(0,1)$. To obtain increasing quantiles we impose, for any $x\in\mathcal{X}$, the function $x^{\prime}\alpha^{0}(u)$ is strictly increasing in $u\in(0,1)$. In forecasting scenarios, $X_{t}$ contains lagged variables. Finally, we note that this model nests the location-scale model, see Section 4.
@@ 2.2 The estimator
with $\|\gamma\|_{1,T}:=\sum_{i=1}^{p}\hat{\sigma}_{i}|\gamma_{i}|$, $\hat{\sigma}_{i}^{2}:=\frac{1}{T}\sum_{t=1}^{T}X_{it}^{2}$.
The second term in the optimization problem in (2) is a penalty term that depends on the magnitude of the elements of the coefficient vector $\gamma$ and the penalty level $\lambda\geq 0$. The norm $\|\cdot\|_{1,T}$ is introduced in Belloni and Chernozhukov, (2011) and used in their quantile regression problem with LASSO penalization for iid data. Belloni et al., (2023) also use it in their LASSO quantile regression for time-series data. It effectively standardizes the data by giving the coefficients corresponding to regressors with high dispersion a larger penalty than coefficients of regressors with low dispersion. We utilize the equivalent norm, such that the LASSO estimator of the quantile coefficients $\alpha^{0}$ uses the same penalization function as the LASSO estimator of the ES coefficients $\gamma^{0}$.
@@ 2.3 Nonasymptotic bound
The nonasymptotic bound on the estimation and prediction errors in Lemma 2.1 below holds on the intersection of the following events: $\mathcal{S}:=\{\max_{i\in[p]}\left|\hat{\sigma}_{i}-\sigma_{i}\right|\leq\frac{1}{4}\}$; $\mathcal{T}:=\{\max_{i\in[p]}\frac{2}{T}\left|\varepsilon^{\prime}X_{j}\right|\leq\lambda_{0}\}$; and $\mathcal{U}:=\{\max_{i,j\in[p]}\left|\frac{1}{T}\sum_{t=1}^{T}X_{it}X_{jt}-E[X_{it}X_{jt}]\right|\leq\lambda_{1}\}$, for positive constants $\lambda_{0}$ and $\lambda_{1}$.
@@ 2.4 Asymptotics
We consider the penalty level
@@ 2.5 Quantile bound
Lemma 2.4 below states that the prediction error $\|\hat{Q}-Q\|_{2}^{2}/T$ converges to zero in probability under Condition QR. Condition QR contains the relevant conditions under which $\hat{\alpha}$ is shown to be consistent in BCMPW. It also imposes a uniform boundedness condition on the regressors, which is sufficient to obtain consistent predictions. We impose it since BCMPW provide a consistency result for the predictions under the alternative norm $\frac{1}{T}\sum_{t=1}^{T}\min\left\{|\hat{Q}_{t}-Q_{t}|,\left(\hat{Q}_{t}-Q_{t}\right)^{2}\right\}$ instead of $\|\hat{Q}-Q\|_{2}^{2}/T$. Lemmas 2.3 and 2.4 together imply $\|\hat{Y}-\tilde{Y}\|_{2}^{2}/T=o_{P}(1)$.
@@
The DGP satisfies the BCMPW conditions as stipulated in Appendix A.3 and the regressors are uniformly bounded, i.e. $\|X_{t}\|_{\infty}\leq M$ (a.s.) for some universal constant $M>0$.
@@ 5 Fuk-Nagaev inequality for beta-mixing
Moreover, there exist uniform constants $C_{q}^{(3)},C_{q}^{(4)}>0$, such that, for any $\delta\in(0,1)$, if $u=C_{q}^{(3)}\frac{(\delta pa_{T})^{1/q}}{d_{T}^{(q-1)/q}}\wedge C_{q}^{(4)}\frac{\sqrt{\log(pa_{T}/\delta)}}{\sqrt{d_{T}}}$, it follows
@@ A.3 Conditions of BCMPW
Consider the DGP in (1). We make the following assumptions.
@@
There exists a $\underline{f}>0$ satisfying
$\inf_{t\in[T],x\in\mathcal{X}}f_{Y_{t}|X_{t}}\left(x^{\prime}\alpha^{0}|x\right)>\underline{f},$
@@
For some $m\geq 0$, we let $\bar{S}(\alpha,m)\in[p]\setminus S_{\alpha,0}$ denote the support of the $m$ largest components in $\alpha$, excluding entries in $S_{\alpha,0}$, denoting the sparsity set of $\alpha^{0}$, and use convention $\bar{S}(\alpha,0)=\emptyset$.
Define the cone the $A^{\alpha}=\{\alpha:\|\alpha_{S_{\alpha,0}^{c}}\|_{1}\leq C_{0}\|\gamma_{S_{\alpha,0}}\|_{1}\}$, for some positive constant $C_{0}$.
@@
For $\alpha\in A^{\alpha}$, let
$J_{\tau}^{1/2}(\alpha):=\sqrt{\frac{\underline{f}}{T}\sum_{t=1}^{T}E\left[\left(X_{t}^{\prime}\alpha\right)^{2}\right]}.$
Then there exists $m\geq 0$ such that
$0<\kappa_{m}:=\inf_{\alpha\in A^{\alpha},\alpha\neq 0}\frac{J_{\tau}^{1/2}(\alpha)}{\left\|\alpha_{S_{\alpha,0}\cup\bar{S}(\alpha,m)}\right\|_{2}}.$
@@
    $q\geq\frac{C}{{\kappa_{0}\underline{f}^{1/2}}}(\phi_{T}\sqrt{a_{T}\log(pa_{T})}\sqrt{s_{0}+1})$, for some sequence $\{\phi_{T}\}$ that satisfies
    $\phi_{T}/(\sqrt{\underline{f}}\log(a_{T}+1))\rightarrow\infty;$
@@
    $\frac{\phi_{T}(1+\sqrt{s_{0}/m})\sqrt{a_{T}\log(pa_{T})}{\sqrt{1+s_{0}}}}{\kappa_{0}\underline{f}^{1/2}\sqrt{d_{T}}}=o(1).$
@@
$\left|\frac{\gamma^{\prime}\hat{\Sigma}\gamma}{\gamma^{\prime}\Sigma\gamma}-1\right|\leq\frac{1}{\gamma^{\prime}\Sigma\gamma}\left|\gamma^{\prime}(\hat{\Sigma}-\Sigma)\gamma\right|\leq\frac{1}{\gamma^{\prime}\Sigma\gamma}\left\|\gamma\right\|_{1}^{2}\|\hat{\Sigma}-\Sigma\|_{\infty}\leq\frac{(1+C_{0})s_{0}}{\phi_{0}^{2}}\lambda_{1}\leq\frac{1}{2},$
@@
$P\left(\max_{i,j\in[p]}\left|\frac{1}{T}\sum_{t=1}^{T}X_{it}X_{jt}-E[X_{it}X_{jt}]\right|>\lambda_{2}\right)$
@@
It can easily be verified that $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|\leq\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$, for any $(y,q,q^{\prime})\in\mathbb{R}^{3}$. Hence,
@@
We have
$\|\hat{Q}-Q\|_{2}^{2}/T=\frac{1}{T}\sum_{t=1}^{T}|X_{t}^{\prime}(\hat{\alpha}-\alpha^{0})|^{2}=\frac{1}{T}\sum_{t=1}^{T}\max_{i\in[p]}|X_{i,t}|^{2}\sum_{i\in[p]}|\hat{\alpha}_{i}-\alpha^{0}_{i}|^{2}\leq M^{2}\|\hat{\alpha}-\alpha^{0}\|_{2}^{2}$
(a.s.), and the result follows from Theorem 4.2 in BCMPW under the imposed conditions.∎
@@
$u$ $=C_{q}^{(3)}\frac{(\delta pa_{T})^{1/q}}{d_{T}^{(q-1)/q}}\wedge C_{q}^{(4)}\frac{\sqrt{\log(pa_{T}/\delta)}}{\sqrt{d_{T}}}$

### v2

@@ 1 Introduction
We use a penalty function that accounts for differences in the scale of the regressors which has not been studied in the literature before for least squares estimators, to our knowledge, and which is borrowed from the penalized quantile regression literature, see Belloni and Chernozhukov, (2011). Although accounting for scale differences is an appealing property on its own, we also prefer this penalty function as it means the penalization function of the coefficients in the VaR and ES estimation steps is equivalent. Indeed, we rely on the quantile regression estimator of Belloni et al., (2023) (from hereon BCMPW) to obtain the VaR predictions in the first step, although our results apply to any quantile estimator that satisfies the conditions.
As financial data is often dependent over time and heavy-tailed, we derive our results under conditions that allow for these properties. Specifically, we allow the data to be $\beta$-mixing sequences with finite moments of a certain order. To obtain our results, we develop a Fuk-Nagaev inequality for $\beta$-mixing sequences, which may be of independent interest. By using the Fuk-Nagaev inequality in conjuction with the same blocking strategy as utilized in BCMPW to obtain approximately independent blocks of data, we obtain rates on the model parameters that are directly comparable to those for the penalized quantile regression estimator in BCMPW. This result is closely related to the Fuk-Nagaev inequality for $\tau$-mixing sequences in Babii et al., (2022), which relies on an alternative blocking strategy, and the tail inequality for sequences that satisfy a sub-Weibull property in Wong et al., (2020).
@@ 2.1 The model
where $\{U_{t}\}$ is a sequence of random errors satisfying $U_{t}|X_{t}\sim\text{Unif}(0,1)$, $\{X_{t}\}$ is a sequence of random vectors satisfying $X_{t}\in\mathcal{X}\subseteq\mathbb{R}^{p}$, and $\alpha^{0}(\cdot)$ is some measurable $p$-dimensional vector functional on $(0,1)$. To obtain increasing quantiles we impose, for any $x\in\mathcal{X}$, the function $x^{\prime}\alpha^{0}(u)$ is strictly increasing in $u\in(0,1)$. In forecasting scenarios, $X_{t}$ contains lagged variables. Finally, we note that this model nests the location-scale model, see Section 4.
@@ 2.2 The estimator
with $\|\gamma\|_{1,T}:=\sum_{i=1}^{p}\hat{\sigma}_{i}|\gamma_{i}|$, $\hat{\sigma}_{i}^{2}:=\frac{1}{T}\sum_{t=1}^{T}X_{it}^{2}$. The second term in the optimization problem in (2) is a penalty term that depends on the magnitude of the elements of the coefficient vector $\gamma$ and the penalty level $\lambda\geq 0$. The norm $\|\cdot\|_{1,T}$ is introduced in Belloni and Chernozhukov, (2011) and used in their quantile regression problem with LASSO penalization for iid data. Belloni et al., (2023) also use it in their LASSO quantile regression for time-series data. It effectively standardizes the data by giving the coefficients corresponding to regressors with high dispersion a larger penalty than coefficients of regressors with low dispersion. We utilize the equivalent norm, such that the LASSO estimator of the quantile coefficients $\alpha^{0}$ uses the same penalization function as the LASSO estimator of the ES coefficients $\gamma^{0}$.
@@ 2.3 Nonasymptotic bound
Lemma 2.1 below provides nonasymptotic bounds on the estimation error $\|\hat{\gamma}-\gamma^{0}\|_{1}$ and the prediction error $\frac{1}{T}\|X(\hat{\gamma}-\gamma^{0})\|_{2}^{2}$ that depend on the sparsity of the ES model and the penalty level, per usual, as well as on the prediction error contained in the feasible auxiliary conditional variables $\hat{Y}$. The bound holds on the intersection of the following events: $\mathcal{S}:=\{\max_{i\in[p]}\left|\hat{\sigma}_{i}-\sigma_{i}\right|\leq\frac{1}{4}\}$; $\mathcal{T}:=\{\max_{i\in[p]}\frac{2}{T}\left|\varepsilon^{\prime}X_{j}\right|\leq\lambda_{0}\}$; and $\mathcal{U}:=\{\max_{i,j\in[p]}\left|\frac{1}{T}\sum_{t=1}^{T}X_{it}X_{jt}-E[X_{it}X_{jt}]\right|\leq\lambda_{1}\}$, for positive constants $\lambda_{0}$ and $\lambda_{1}$.
@@ 2.4 Asymptotics
Lemma 2.2 below gives a consistency result on the estimation and prediction errors, for penalty levels that satisify
@@ 2.5 Quantile bound
Lemma 2.4 below states that the prediction error $\|\hat{Q}-Q\|_{2}^{2}/T$ converges to zero in probability under Condition QR. Condition QR contains the relevant conditions under which $\hat{\alpha}$ is shown to be consistent in BCMPW. Lemmas 2.3 and 2.4 together imply $\|\hat{Y}-\tilde{Y}\|_{2}^{2}/T=o_{P}(1)$.
@@
The DGP satisfies the BCMPW conditions as stipulated in Appendix A.3.
@@ 5 Fuk-Nagaev inequality for beta-mixing
Moreover, there exist uniform constants $C_{q}^{(3)},C_{q}^{(4)}>0$, such that, for any $\delta\in(0,1)$, if $u=C_{q}^{(3)}\frac{(pa_{T}/\delta)^{1/q}}{d_{T}^{(q-1)/q}}\wedge C_{q}^{(4)}\frac{\sqrt{\log(pa_{T}/\delta)}}{\sqrt{d_{T}}}$, it follows
@@ A.3 Conditions of BCMPW
Consider the DGP in (1). We impose the following assumptions.
@@
There exists a $\underline{f}\in(0,1]$ satisfying
$\inf_{x\in\mathcal{X}}f_{Y_{t}|X_{t}}\left(x^{\prime}\alpha^{0}|x\right)>\underline{f},$
@@
As an example of a conditional distribution of $Y_{t}$ given $X_{t}$, consider the location-scale model $Y_{t}=E[Y_{t}|X_{t}]+\text{std}(Y_{t}|X_{t})\nu_{t}$, where $E[Y_{t}|X_{t}]=X_{t}^{\prime}\xi$ and $\text{std}(Y_{t}|X_{t})=X_{t}^{\prime}\zeta$, with location and scale parameter vectors $\xi\in\mathbb{R}^{p}$ and $\zeta\in\mathbb{R}^{p}$, respectively. We impose positive standard deviations and independent errors. Specifically, (i) $\inf_{x\in\mathcal{X}}x^{\prime}\zeta>0$; (ii) $\nu_{t}\sim iid\,\,F_{\nu}$ for some distribution $F_{\nu}$ with continuously differentiable cdf and $\nu_{t}$ independent of $X_{t}$ and its past. Clearly, this model can be expressed in terms of the model in Equation (1) as follows: $Y_{t}=X_{t}^{\prime}\alpha^{0}(U_{t})$, where $\alpha^{0}(u)=\xi+F_{\nu}^{-1}(u)\zeta$, with $F_{\nu}^{-1}(\cdot)$ denoting the inverse cdf. It follows the conditional pdf can be expressed as $f_{Y_{t}|X_{t}}(y|x)=\frac{1}{x^{\prime}\zeta}f_{\nu}\left((y-x^{\prime}\xi)/(x^{\prime}\zeta)\right)$.
We let $S_{\alpha,0}$ denote the sparsity set of $\alpha^{0}$, and let $s_{\alpha,0}$ denote the cardinality of $S_{\alpha,0}$. Here, we specifically allow for different sparsity sets for the quantile and ES model parameters, respectively, $\alpha^{0}$ and $\gamma^{0}$. Moreover, we define the cone $A^{\alpha}=\{\alpha:\|\alpha_{S_{\alpha,0}^{c}}\|_{1}\leq C_{0}\|\alpha_{S_{\alpha,0}}\|_{1}\}$, for some positive constant $C_{0}$.
@@
The following conditions are satisfied.
  1. 1.
    For any $\alpha\in A^{\alpha}$, let
    $J_{\tau}(\alpha):=\underline{f}E\left[\left(X_{t}^{\prime}\alpha\right)^{2}\right]=\underline{f}\alpha^{\prime}\Sigma\alpha.$
    It holds
    $0<\kappa_{0}:=\inf_{\alpha\in A^{\alpha},\alpha\neq 0}\frac{J_{\tau}^{1/2}(\alpha)}{\left\|\alpha_{S_{\alpha,0}}\right\|_{2}};$
  2. 2.
@@
    $q\geq\frac{C}{{\kappa_{0}\underline{f}^{1/2}}}(\psi_{T}\sqrt{a_{T}\log(pa_{T})}\sqrt{s_{\alpha,0}+1})$, for some universal constant $C>0$ and any sequence $\{\psi_{T}\}$ that satisfies
    $\psi_{T}/\left(\sqrt{\underline{f}}\log(a_{T}+1)\right)\rightarrow\infty;$
@@
    $\frac{\psi_{T}\sqrt{(1+s_{\alpha,0})a_{T}\log(pa_{T})}}{\sqrt{d_{T}}\kappa_{0}\underline{f}}=o(1)$;
  3. 3.
    $\frac{\underline{f}s_{\alpha,0}}{\kappa_{0}^{2}}\frac{(pa_{T})^{2/q}}{d_{T}^{(q-1)/q}}=o(1)$.
We note that Conditions B2.1 and B3 are closely related to REC and R. Indeed, Condition B2.1 can be rewritten as $\|\alpha_{S_{\alpha,0}}\|_{1}^{2}\leq s_{\alpha,0}\|\alpha_{S_{\alpha,0}}\|_{2}^{2}\leq\frac{s_{\alpha,0}\underline{f}}{\kappa_{0}^{2}}\alpha^{\prime}\Sigma\alpha$, for any $\alpha\in A_{\alpha}$, which is similar to Condition REC. Moreover, if we assume $\kappa_{0}$ and $\underline{f}$ are universal constants (like $\phi_{0}$ in REC), then the rates in B3.2 and B3.3 no longer depend on $\kappa_{0}$ and $\underline{f}$, and are similar to those in Condition R (ignoring $\psi_{T}$). Condition B2.2 represents the restricted nonlinear impact condition introduced in Belloni and Chernozhukov, (2011) and is standard in the $\ell_{1}$-penalized quantile regression literature.
@@
$\left|\frac{\gamma^{\prime}\hat{\Sigma}\gamma}{\gamma^{\prime}\Sigma\gamma}-1\right|\leq\frac{1}{\gamma^{\prime}\Sigma\gamma}\left|\gamma^{\prime}(\hat{\Sigma}-\Sigma)\gamma\right|\leq\frac{1}{\gamma^{\prime}\Sigma\gamma}\left\|\gamma\right\|_{1}^{2}\|\hat{\Sigma}-\Sigma\|_{\infty}\leq\frac{(1+C_{0})^{2}s_{0}}{\phi_{0}^{2}}\lambda_{1}\leq\frac{1}{2},$
@@
$P\left(\max_{i,j\in[p]}\left|\frac{1}{T}\sum_{t=1}^{T}X_{it}X_{jt}-E[X_{it}X_{jt}]\right|>\lambda_{1}\right)$
@@
We may use the inequality $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|\leq\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$, which holds for each $(y,q,q^{\prime})\in\mathbb{R}^{3}$. Hence, we obtain the result
@@
Finally, we verify the inequality $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|\leq\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$, for each $(y,q,q^{\prime})\in\mathbb{R}^{3}$ by considering four cases that encompass all orderings of $(y,q,q^{\prime})$. Case 1: If $y<q$ and $y<q^{\prime}$, then $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|=|q-q^{\prime}-\frac{1}{\tau}(q-q^{\prime})|\leq\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$; Case 2: If $q<y\leq q^{\prime}$, then $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|=|q-q^{\prime}+\frac{1}{\tau}(y-q^{\prime})|\leq|q-q^{\prime}|+\frac{1}{\tau}|y-q^{\prime}|\leq\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$, where the first inequality follows by the Triangle Inequality and the second inequality from the ordering $q<y<q^{\prime}$; Case 3: If $q<y\leq q^{\prime}$, then $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|=|q-q^{\prime}-\frac{1}{\tau}(y-q)|\leq|\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$, by similar steps as in Case 2; Case 4: If $q\leq y$ and $q^{\prime}\leq y$, then then $|\xi_{\tau}(y,q)-\xi_{\tau}(y,q^{\prime})|=|q-q^{\prime}|\leq|\left(1+\frac{1}{\tau}\right)|q-q^{\prime}|$.
@@
Under the imposed conditions, the Assumptions of Theorem 4.2 in BCMPW are satisfied, which is shown at the end of this proof.
Then, notice
$\|\hat{Q}-Q\|_{2}^{2}/T$ $=\frac{1}{T}\sum_{t=1}^{T}\left(X_{t}^{\prime}(\hat{\alpha}-\alpha^{0})\right)^{2}$
$=(\hat{\alpha}-\alpha^{0})^{\prime}\frac{1}{T}\sum_{t=1}^{T}X_{t}X_{t}^{\prime}(\hat{\alpha}-\alpha^{0})$
$=(\hat{\alpha}-\alpha^{0})^{\prime}\hat{\Sigma}(\hat{\alpha}-\alpha^{0}).$
We may employ the inequality $\alpha^{\prime}\hat{\Sigma}\alpha\leq C_{1}\alpha^{\prime}\Sigma\alpha$, for any $\alpha\in A_{\alpha}$, which holds for some universal constant $C_{1}>0$ with probability approaching one, as shown at the end of this proof. It follows that we may bound the RHS (with probability approaching one) by
$C_{1}(\hat{\alpha}-\alpha^{0})^{\prime}\Sigma(\hat{\alpha}-\alpha^{0})=\frac{C_{1}}{\underline{f}}J_{\tau}(\hat{\alpha}-\alpha^{0}),$
since $\hat{\alpha}-\alpha^{0}\in A_{\alpha}$ with probability approaching one by Lemma S4 in BCMPW and where we recall the definition of $J_{\tau}(\cdot)$ in Section A.3.
Finally, in the proof to their Theorem 4.2, BCMPW establish
$J_{\tau}(\hat{\alpha}-\alpha^{0})^{1/2}\lesssim\frac{\psi_{T}\sqrt{(1+s_{\alpha,0})a_{T}\log(pa_{T})}}{\sqrt{d_{T}}\kappa_{0}\underline{f}^{1/2}},$
with probability approaching one, such that
$\|\hat{Q}-Q\|_{2}^{2}/T=O_{P}\left(\frac{\psi_{T}^{2}(1+s_{\alpha,0})a_{T}\log(pa_{T})}{d_{T}\kappa_{0}^{2}\underline{f}^{2}}\right),$
and hence $\|\hat{Q}-Q\|_{2}^{2}/T=o_{P}(1)$ under the rate in Condition QR.
We now establish the inequality $\alpha^{\prime}\hat{\Sigma}\alpha\leq C_{1}\alpha^{\prime}\Sigma\alpha$, for any $\alpha\in A_{\alpha}$, holds for some universal constant $C_{1}>0$ with probability approaching one. As in the proof of Lemma 2.1, for any $\alpha\in A_{\alpha}$, under the Condition QR and on $\mathcal{U}$,
$\left|\frac{\alpha^{\prime}\hat{\Sigma}\alpha}{\alpha^{\prime}\Sigma\alpha}-1\right|\leq\frac{1}{\alpha^{\prime}\Sigma\alpha}\left|\alpha^{\prime}(\hat{\Sigma}-\Sigma)\alpha\right|\leq\frac{1}{\alpha^{\prime}\Sigma\alpha}\left\|\alpha\right\|_{1}^{2}\|\hat{\Sigma}-\Sigma\|_{\infty}\leq\frac{(1+C_{0})^{2}\underline{f}s_{0}}{\kappa_{0}^{2}}\lambda_{1}\leq\frac{1}{2}\underline{f}\leq\frac{1}{2}.$
such that $\gamma^{\prime}\Sigma\gamma\leq 2\gamma^{\prime}\hat{\Sigma}\gamma$ with probability approaching one if $P(\mathcal{U})\rightarrow 1$ for $\lambda_{1}\leq\frac{\kappa_{0}^{2}}{2(1+C_{0})^{2}\underline{f}s_{0}}$. From similar steps to those in the proof of Lemma 2.2, we obtain
$P(\mathcal{U})\leq 3p^{2}a_{T}\Biggl{(}\frac{C_{q}^{(1)}\underline{f}^{q}s_{0}^{q}}{\kappa_{0}^{2q}d_{T}^{q-1}}+\exp\bigl{(}-C_{q}^{(2)}\kappa_{0}^{4}d_{T}/(s_{0}^{2}\underline{f}^{2})\bigr{)}\Biggr{)}+2p^{2}d_{T}\bar{\beta}(a_{T}),$
such that the RHS converges to zero under the rates imposed in Condition QR.
Finally, we show the Assumptions 2-4 in BCMPW imposed in their Theorem 4.2 are satisfied. Assumption 2 is satisfied under Conditions A, D, MOM, and R; see Lemma 2.2. Assumption 3 is satified since we can decompose $Y_{t}$ as $Y_{t}=X_{t}^{\prime}\alpha^{0}(\tau)+X_{t}^{\prime}\left(\alpha^{0}(U_{t})-\alpha^{0}(\tau)\right)$. Defining $\epsilon_{t}=X_{t}^{\prime}\left(\alpha^{0}(U_{t})-\alpha^{0}(\tau)\right)$ it is obvious that $P(\epsilon_{t}\leq 0|X_{t})=\tau$ from the definition of the vector functional $\alpha^{0}(\cdot)$. Indeed, by monotonicity of $X_{t}^{\prime}\alpha^{0}(\tau)$ in $\tau$ (a.s.) we obtain,
$P(\epsilon_{t}\leq 0|X_{t})$ $=P\left(X_{t}^{\prime}\alpha^{0}(U_{t})\leq X_{t}^{\prime}\alpha^{0}(\tau)|X_{t}\right)=P\left(U_{t}\leq\tau\right)=\tau.$
Moreover, given the measurability of $\alpha^{0}(\cdot)$ and the $U_{t}$ being iid, it follows $\{\alpha^{0}(U_{t})\}$ is mixing of any rate and therefore $\{Y_{t}\}$ and $\{\epsilon_{t}\}$ are mixing of rate $\overline{\beta}(t)$ by Theorem 3.49 in White, (2001). Hence, by another application of Theorem 3.49, $\{(X_{t},\epsilon_{t})\}$ is stationary and $\beta$-mixing of the same rate. Finally, the requirements on the conditional density are reproduced in Condition QR in an equivalent form for strictly stationary data. Assumption 4 is reproduced in Condition QR for $m=0$. The bound on $q$ in the statement of Theorem 4.2 is also included. ∎
@@
$u$ $=C_{q}^{(3)}\frac{(pa_{T}\delta)^{1/q}}{d_{T}^{(q-1)/q}}\wedge C_{q}^{(4)}\frac{\sqrt{\log(pa_{T}/\delta)}}{\sqrt{d_{T}}}$